parrallel graph of 2x -y =1 Y intercept (o,-3)
Answer by josgarithmetic(39617)

parrallel graph of 2x -y =1 Y intercept (o,-3)

Imagining you really mean,line parrallel to graph of 2x -y =1 and with a Y intercept (0,-3)
(Number 0 is better than the letter o when you really do mean "zero").

You want the equation 2x-y=c which will give the same slope as 2x-y=1 and therefore will be parallel; but we do not yet know c. Plug in the given y-intercept and find c:

2*0-(-3)=c
c=3
Your wanted line is 2x-y=3